OpenStudy (xochital):

Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression.

OpenStudy (anonymous):

Is this the equation you are describing? \[3\sqrt{2x} + x \sqrt{8x} - 5\sqrt{18x} = -10x \sqrt{2x}\]

OpenStudy (xochital):

yes :)

OpenStudy (anonymous):

The first step would be to simplify all the square roots. For example: \[\sqrt{90} = \sqrt{9} * \sqrt{10} = 3 \sqrt{10}\] 9 is the only perfect square factor of 90. Does that make sense?

OpenStudy (xochital):

yes, it does

OpenStudy (anonymous):

Can you do that with each of the terms in your question?

OpenStudy (xochital):

i don't think so... aren't they already simplified?

OpenStudy (anonymous):

No. Factor each of coefficients under the radical to find which ones have perfect squares. coefficient = number multiplied by the variable radical = square root sign

OpenStudy (xochital):

oh okay, so simplified it would be... 3*sqrt(2x) + 2x*sqrt(2x) - 15*sqrt(2x)

OpenStudy (anonymous):

Exactly! Now, combine like terms.

OpenStudy (xochital):

sqrt(2x) * (3 + 2x - 15) which equals sqrt(2x) * (-12 + 2x)... so nakim must have combined 3 + 2x - 15 to get -10x.

OpenStudy (anonymous):

Yes.

OpenStudy (xochital):

oh okay thanks, i understand it better now :)

OpenStudy (anonymous):

Great!